OpenStudy (anonymous):

the probability that a professional baseball player will get a hit is 1/3. Calculate the exact probability that he will get at least 3 hits in 5 attempts.

OpenStudy (anonymous):

that was wrong, sorry

OpenStudy (anonymous):

at least 3 means 3 or 4 or 5 you have to compute each of these and add them up

OpenStudy (anonymous):

exactly 3 is \[P(x=3)=\binom{5}{3}\left(\frac{1}{3}\right)^3\left(\frac{2}{3}\right)^2\] meaning 3 hits, 2 not hits, and \(\binom{5}{3}=\frac{5\times 4}{2}=10\) ways to get order the three hits and two not hits

OpenStudy (anonymous):

The answer has to be the exact probability so shouldnt it be written differently?

OpenStudy (anonymous):

yeah you have to actually compute that number, and then do it again for \(P(x=4)\) and \(P(x=5)\)

OpenStudy (anonymous):

\[P(x=4)=\binom{5}{4}\left(\frac{1}{3}\right)^4\left(\frac{2}{3}\right)\]and \[P(x=5)=\left(\frac{1}{3}\right)^5\]

OpenStudy (anonymous):

For the answer i got 17/81 correct?